OpenStudy (anonymous):

Find the absolute extrema for these functions: f(x) = x^3 - 3x^2 for [-1,3]

OpenStudy (anonymous):

wouldn't max be 2 not 0?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

first take the derivative and set it equal to 0 to find the critical points. so: f '(x) = 3x^2 - 6x = 0 0 = x(3x -6) x=0 , x=2 "candidates for extrema are critical points and endpoints so.... f(0) =0 f(2) = -4 f(-1) = -4 f(3) = 0

OpenStudy (anonymous):

ok but wouldn't maximum be 2 not 0?

OpenStudy (anonymous):

oh wait nvm lmfao

OpenStudy (anonymous):

@TweT226 can u help me with another one pls? ^^

OpenStudy (anonymous):

yeah?